# Do Androids Know They’re Only Dreaming of Electric Sheep?

### v1

@@ 6 Discussion
Our work is motivated by the need for efficient LLM hallucination evaluation. Though computationally efficient, one of the key limitations of probing is in the need for labeled in-domain data for probe training. Our results in Table˜8 suggest poor generalization to out-of-domain tasks, although perhaps this would be mitigated by larger and more diverse training sets. Furthermore, it perhaps goes without saying that probing requires access to the hidden states themselves. If LLMs continue to move behind closed-source APIs, such hidden state probes will not be possible for those without privileged access.
Test Train F1-R
CDM CF 0.44
E2E 0.44
CF+E2E 0.46
CDM 0.45
CF CDM 0.60
E2E 0.80
CDM+E2E 0.75
CF 0.88
E2E CDM 0.60
CF 0.75
CDM+CF 0.76
E2E 0.87
Table 8: Response-level PoolingE cross-task probe performance, as measured by F1-R. All experiments use the same amount of training data (a random subset of the full training set). The "+" sign indicates an equal mixture of two training tasks.
Ecological Validity.
The conclusions from an ecologically valid experiment should hold outside the context of that study De Vries et al. (2020). Synthetic hallucinations are generally ecologically invalid for organic hallucination detection. Extending beyond surface-level error distributions Goyal and Durrett (2021), these differences in internal representations further suggest that language model behavior during hallucination fundamentally misaligns with current expectations of ungrounded generations. Training on the same modality—synthetic or organic—is imperative to maximize hallucination detection performance, and our results on the synthetic modality, especially on CNN / DailyMail, suggest probing to be a feasible method for complex multi-sentence entailment classification that can leverage the capability of LLMs.
@@
Despite asking annotators to label what they deem to be the minimal spans of hallucination in a response, considerable variation still exists between annotators on judgments of their exact span boundaries. It remains an open questions how to improve annotator agreement in this context.
@@
Even if span boundaries were standardized, they may not align perfectly with hallucination signals in LLM internal representations. For instance, while a generation may first contradict a given knowledge source at a certain token, the ultimate signal that it will hallucinate may already be salient in the tokens preceding it. Alternatively, the signal might be computed only later, as the decoder encodes previously generated tokens and recognizes that it is currently pursuing an ungrounded train of thought. Thus, there may be probe architectures that could improve hallucination detection performance and mitigation. Nonlinear probes such as feed-forward networks are also worth considering: the hidden layer(s) of a feed-forward network could learn to detect different types of hallucinations.
@@ Acknowledgments
We thank Val Ramirez, Patrick Xia, Harsh Jhamtani, Hao Fang, Tongfei Chen, and Radhika Gaonkar for their helpful comments, thoughts, and discussions, as well as the data specialists who contributed and aided in the creation of this work.
@@
Here, we provide additional information about annotation guidelines, annotator disagreement examples, prompts used to generate organic responses to our tasks, and ChatGPT baseline comparisons.
@@ Appendix B Annotator Disagreements
Examples of annotator disagreements in the exact spans of hallucination are shown in Table˜9. Qualitatively, we find that annotators agreed generally on the specific items of hallucination; i.e. the subject that was hallucinated (i.e. the taxi driver), the description of what was mentioned (i.e. the description of accounting), and the attribute that was hallucinated (i.e. the specific rating of the resturant). However, where exactly those hallucinated spans began and ended was a source of reasonable variation among annotators.
Appendix C Prompts
@@ C.1 ChatGPT Prompts
For the most basic zero-shot prompt classifier with no reasoning, we use the following prompt for CNN / DailyMail, following Luo et al. (2023):
Decide if the following summary is consistent with the corresponding article. Note that consistency means all information in the summary is supported by the article.Article: [Article]Summary: [Summary]Answer (yes or no):

### v2

@@ 7 Conclusion
Our work is motivated by the need for efficient LLM hallucination evaluation. Though computationally efficient, one of the key limitations of probing is in the need for labeled in-domain data for probe training. Our results in Table 8 suggest poor generalization to out-of-domain tasks, although perhaps this would be mitigated by larger and more diverse training sets. Furthermore, it perhaps goes without saying that probing requires access to the hidden states themselves. If LLMs continue to move behind closed-source APIs, such hidden state probes will not be possible for those without privileged access.
@@
Labeling for faithfulness errors remains a challenging task for annotators Ladhak (2024). Despite asking annotators to label what they deem to be the minimal spans of hallucination in a response and providing a considerably detailed annotation schema (Appendix A), sizable variation still exists between annotators on judgments of their exact span boundaries. It remains an open question how to reduce or circumvent annotator disagreement in this context.
@@
Even if span boundaries were standardized, they may not align perfectly with hallucination signals in LLM internal representations. For instance, while a generation may first contradict a given knowledge source at a certain token, the ultimate signal that it will hallucinate may already be salient in the tokens preceding it. Alternatively, the signal might be computed only later, as the decoder encodes previously generated tokens and recognizes that it is currently pursuing an ungrounded train of thought. Thus, there may be probe architectures that could improve hallucination detection performance and mitigation. Nonlinear probes such as feed-forward networks are also worth considering: the hidden layer(s) of a feed-forward network could learn to detect different types of hallucinations. Furthermore, the lack of task transfer that we currently observe presents interesting future challenges in increasing the generalizability of our approach in search of a more consistent signal of hallucination, if one exists.
@@ Acknowledgments
We thank Val Ramirez, Patrick Xia, Harsh Jhamtani, Hao Fang, Tongfei Chen, Radhika Gaonkar, and Tatsunori Hashimoto for their helpful comments, thoughts, and discussions, as well as the data specialists who contributed and aided in the creation of this work.
@@
Here, we provide additional information about annotation guidelines, annotator disagreement examples, prompts used to generate organic responses to our tasks, ChatGPT baseline comparisons, and expanded results.
@@ Appendix B Annotator Disagreements
Examples of annotator disagreements in the exact spans of hallucination are shown in Table 7. Qualitatively, we find that annotators agreed generally on the specific items of hallucination; i.e. the subject that was hallucinated (i.e. the taxi driver), the description of what was mentioned (i.e. the description of accounting), and the attribute that was hallucinated (i.e. the specific rating of the resturant). However, where exactly those hallucinated spans began and ended was a source of reasonable variation among annotators.
Appendix C Prompts and Further Evaluation Details
@@
FActScore Evaluation.
As FActScore Min et al. (2023) was originally designed for evaluating the faithfulness of long-form generated texts, we slightly modify its evaluation for our Data-to-Text task, E2E. Specifically, instead of evaluating on raw attribute sets (i.e. name[Clowns], eatType[pub], near[The Sorrento]), we transform these attributes into natural language sentences via simple regular expressions. In this example, these attributes would instead be transformed into:
The name of this restaurant is Clowns. This restaurant is a pub. This restaurant is near The Sorrento.
For our other tasks on abstractive summarization and knowledge-grounded dialogue generation, we keep evaluations unchanged with the original implementation of FActScore for our evaluations.
@@ C.1 ChatGPT Prompts
For the most basic zero-shot prompt classifier with no reasoning, we use the following prompt for CNN / DailyMail, making minimal changes relative to Luo et al. (2023):
Decide if the following summary (SUMMARY) is consistent with the corresponding article (ARTICLE). Note that consistency means all information in the summary (SUMMARY) is supported by the article (ARTICLE). Article (ARTICLE): [Article] Summary (SUMMARY): [Summary] Answer (yes or no):
@@
Appendix D Expanded Results and Discussions
Grounding Behavior Saliency.
Full saliency results for all tasks and layers are shown in Figure 4; saliency as stratified across hallucination layers and types across both MLP and Attention probes is shown in Figure 5.
Refer to captionRefer to captionRefer to caption
Figure 4: Grounding behavior saliency—as measured by probe response-level F1 on test set—of single layer probe classifiers trained on the attention-head out projections and feed-forward states of llama-2-13b, stratified across tasks and layers. Vertical dashed lines denote the layers that respective probes first surpass 95% of their peak performance.
Figure 5: Intrinsic and extrinsic hallucination saliency across layers for synthetic abstractive summarization (BUMP, for CDM) as measured by response-level MLP and Attention probe F1 on hallucination types.
Task Transfer.
Full results for task transfer using the response-level PoolingE probe are shown in Table 8.
Test Train F1-R
CDM CF 0.67
E2E 0.65
CF+E2E 0.68
CDM 0.70
CF CDM 0.84
E2E 0.81
CDM+E2E 0.84
CF 0.93
E2E CDM 0.06
CF 0.71
CDM+CF 0.71
E2E 0.86
Table 8: Response-level PoolingE cross-task probe performance, as measured by F1-R. All experiments here use the same amount of training data (a random subset of the full training set). The "+" sign indicates an equal mixture of two training tasks.
Differences from Specific Prior Work.
In the review process, reviewers raised uncertainties about the specific differences between our work and that of Azaria and Mitchell (2023) and Kadavath et al. (2022); we note our response here.
Azaria and Mitchell (2023) are interested in detecting the truth or falsehood of factual statements (e.g., “H2O is water, which is essential for humans”). The models they probe are presented with these statements without any grounding knowledge, so the probe can only measure if the model’s hidden states reveal some discrepancy between the encoded statement and the training data as encoded by the learned weights. This cannot work unless the learned weights really encode the truth, so one should only test facts that are well-known and unchanging or slowly changing in time, modality, etc.
In contrast, a RAG chatbot built on a company’s internal data will be discussing specialized and often ephemeral retrieved information that is not well represented in the world knowledge of an LLMs weights. Here, it is important to ensure that the model appropriately follows task instructions, and in particular that it does not hallucinate in the sense defined by § 1 and Ji et al. (2023): its responses should be appropriately grounded in the retrieved data.
For emphasis we highlight three important distinctions. (1) Azaria and Mitchell (2023) examine the veracity of statements without supporting text, while we examine hallucination with respect to grounding text and task instructions. (2) They do not probe models on text that the models themselves generate, but text provided by a secondary model (ChatGPT) and additional human curation. This is an important distinction—we evaluate hallucination on both organically generated texts (and their hidden states) and those generated by a secondary model + human manual curation (synthetic grounding errors) and find that the probe performance generally degrades when probing across these two cases. Finally, (3) they do not study span or token-level detection, which is one of our primary contributions.
Kadavath et al. (2022) similarly probe an LLM in a closed book QA task by training an attention head to detect when the model does not know the answer to a trivia question in the TriviaQA dataset Joshi et al. (2017). This is different than detecting hallucination in a grounded generation task, as it only tests recall of knowledge already stored in model weights. As for Azaria and Mitchell (2023), their generation task does not require the model to draw information from evidence in the prompt that may disagree with stored knowledge.